Let P(x) = x 2 − x − 6. Compute P(A) for A =  3 −1 0 −2 


1
Expert's answer
2021-05-12T02:08:21-0400

Given,

"P(x)=x^2-x-6"

"=x^2-3x+2x-6"

"=x(x-3)+2(x-3)"

"=(x-3)(x+2)"

As here matrix "A=\begin{bmatrix}
3 & -1\\
0 &2 
\end{bmatrix}"

Matrix order is "2\times 2"

Hence, "A^2=\begin{bmatrix}
3 & -1\\
0 & 2 
\end{bmatrix}\begin{bmatrix}
3 & -1\\
0 & 2 
\end{bmatrix}"


"=\begin{bmatrix}
9 & (-3-2)\\
0 & 4 
\end{bmatrix}"


"=\begin{bmatrix}
9 & (-5)\\
0 & 4 
\end{bmatrix}"

Now, substituting the values,

"P(A)=A^2-A-6"

Hence,

"=\begin{bmatrix}
9 & (-5)\\
0 & 4 
\end{bmatrix}-\begin{bmatrix}
3 & -1\\
0 &2 
\end{bmatrix}-6\begin{bmatrix}
1 & 0\\
0 &1 
\end{bmatrix}"


"=\begin{bmatrix}
0 & -4\\
0 &0
\end{bmatrix}"


Hence P(A)"=\begin{bmatrix}
0 & -4\\
0 &0
\end{bmatrix}" .